OpenStudy (anonymous):

if you substitute x= -1/2 on this equation 2x^2=7x+4, do you get the same answer on both side?

OpenStudy (anonymous):

my bad on the last equation, sorry.

OpenStudy (anonymous):

ok 2x^2 = 1/2 as before 7(-1/2) +4 = -3 1/2 + 4 = 1/2 yes